[Hallc_running] [New Logentry] Hall C RC daily update

simona at jlab.org simona at jlab.org
Wed Aug 5 15:00:03 EDT 2020

Logentry Text:
Hall C did not get any beam since the storm down. There was hope that we may get CW beam by the end of the owl shift today but when MCC started to bring C back online after 6 am a problem with beam non-collinearity at the injector was observed: https://logbooks.jlab.org/entry/3820807. The injector experts are still working to resolve it: https://logbooks.jlab.org/entry/3820876 , https://logbooks.jlab.org/entry/3820888 , https://logbooks.jlab.org/entry/3820927

Hall B will need about two more shifts of data taking before they go to 5th pass (which will allow A to take 1 pass beam). The plan as discussed at the scheduling meeting today and as outlined on the white board is to conclude the 1 pass running in B and do their pass change, then solidify the 5th pass machine by ramping up CW to Hall C to stable 30 muA and then ramp up CW to A (1 pass).

Our counting house is fully staffed waiting for beam. This morning the target polarization looked great: https://logbooks.jlab.org/entry/3820866
Once beam delivery to C is resumed we will do harp scans with the 3H07A and B to check the beam profile and if it looks acceptable we will start production running.

This is my last update as Hall C RC for this running period. Holly is taking over as your RC today.

We will have an RC meeting at 4:15 pm. This is how to connect via bluejeans:  https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3821006
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200805/568bba58/attachment.html>

More information about the Hallc_running mailing list